QUESTION 4 1 points Save Answer You rub an inflated balloon on the carpet and it acquires a potential that is 1,564 V lower than its potential before it became charged. If the charge is uniformly distributed over the surface of the balloon and if the radius of the balloon is 15.7 cm, what is the net charge on the balloon? (Give your answer in decimal using nC as unit
Answer #1

Initially the balloon is uncharged. So that the potential is zero. When it is rubbed, its potential becomes -1.564V. That is, the potential at the surface of the balloon is -1.564V since it is charged uniformly. Assuming that the balloon is a perfect sphere of radius R. The potential at the surface of a charged sphere of radius R is given by,

V=\frac{Q}{4\pi\epsilon_0R}

where Q is the total charge of the sphere.

Then, the charge of the balloon is,

Q=4\pi\epsilon_0RV=4\times\pi\times 8.854\times10^{-12}\times (-1.564) \ C\\ Q=-2.7320\times 10^{-11} \ C

or,

Q=-0.0273 \ nC

Please don't forget to give thumbs up to the answer. Best wishes!